OpenStudy (anonymous):

passes through (0,1) and (2,5)

hero (hero):

What passes through those points?

OpenStudy (anonymous):

all i know is that m=slope

hero (hero):

So you need to find m

OpenStudy (anonymous):

thats what im tryna do and i cnt

hero (hero):

If y = mx + b and you have to points (x,y) = (0,1) = (2,5), then plug each point into the line formula: 5 = 2m + b 1 = 0m + b Next isolate b in both: 5 - 2m = b 1 = b Now substitute b = 1 into the first equation: 5 - 2m = 1 Now isolate m: 5 - 1 = 2m 4 = 2m 4/2 = m 2 = m Therefore, the slope is m = 2

hero (hero):

Do you have another one?